OpenStudy (anonymous):

Which of the following is the LCM of the polynomials? (x2 + 10x + 9) and (x2 - 1) A.(x + 1) B.(x + 1)( x + 9) C.(x + 1)(x - 1) D.(x + 1)(x - 1)(x + 9)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

well, you have to factor both to find out: x^2+10x+9=(x+1)(x+9) x^2-1=(x+1)(x-1) Thus we can easily see that your LCM is (x+1)